# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

### v2

@@ 3.2 Few-Shot Generalization to In-Distribution Data
Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:N_{M},t=0:N}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration. In the in-distribution (IID) setting, for each test task $M^{j}$, we use behavior policies only from the same tasks $\{\pi_{\beta}^{j,t}\}_{j,t=0:N}$ to collect the context and the conditioned rollouts, as visualized in Figure 1. Figure 4 illustrates the learning curves of UNICORN vs. all baselines, which correspond to the IID entries of Table 2. The results demonstrate that UNICORN variants, especially UNICORN-SS, consistently achieve SOTA performance across all benchmarks. The observation that the asymptotic performance of CSRO is comparable to UNICORN is expected since it also optimizes a linear combination of the lower and upper bound of $I(\bm{Z};\bm{M})$ by Theorem 2.3, with a different implementation choice. Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process. Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired. Moreover, the average training time for MACAW is almost three times longer than the context-based methods.
@@ 3.3 Few-Shot Generalization to Out-of-Distribution Behavior Policies
To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:N_{M},t=0:N}}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1. The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 2. While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin. Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT. This also delineates the storyline of COMRL development from 2021 to date
@@ 3.4 Influence of Data Quality
Table 3: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
@@
We observe that CORRO fails on all three datasets. This may be due to the reliance of CORRO on the negative sample generator. Under narrow data distribution or poor random data distribution, it might be more challenging to generate negative samples of good quality. As for Prompt-DT, its performance improves significantly with increased data quality, which is expected since the decision transformer adopts a behavior-cloning-like supervised learning style.
@@ 4.1 Is UNICORN Model-Agnostic?
As UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic. Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT. Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$. We name these variants as UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
As shown in Table 4 and Figure 7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option. Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [43] and DT [44]. We leave this verification to future work.
@@ 4.2 Can UNICORN be Exploited for Model-Based Paradigms?
Figure 5: Testing returns for OOD tasks. The learning curve is averaged by 6 random seeds.
With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$, UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks. We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1). Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$. The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent. We adopt the conventional ensemble technique in model-based RL  [45, 46] to stabilize the training process.
As shown in Figure 5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
@@ 5 Conclusion & Limitation
In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective. Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift. Based on the proposed framework, we instantiate two novel algorithms called UNICORN-SUP and UNICORN-SS which are demonstrated to be remarkably robust and broadly applicable through extensive experiments. We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL paradigms. Mathematically, we expect our framework to be able to incorporate almost all existing COMRL methods that focus on task representation learning. Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic of w.r.t , it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound like the ones introduced by Theorem 2.3, up to some regularizations or constraints.
Limitation Since our framework assumes a decoupling of task representation learning and offline policy optimization, it does not directly elucidate how high-quality representations are indicative of higher downstream RL performance. Another limitation would be the scale of the experiments. Our offline datasets cover at most 40 tasks and $\sim 450k$ transitions, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small in Theorem 2.4). However, we expect our conclusion to extrapolate to larger task sets, dataset sizes and model parameters, for which the DT variants may demonstrate better scaling. Lastly, in principle, the information theoretic formalization in this work should be directly applicable to online RL. However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B), which are apparently violated in the online scenario. Extending our framework to online RL is interesting and nontrivial, which we leave for future work.
@@ Appendix B Proofs
Remark: Consider a practical implementation of $p(\bm{z}|M)$ as a multivariate Gaussian $\mathcal{N}(\bm{\mu}_{M},\bm{\Sigma}_{M})$, which gives $H(\bm{Z}|\bm{M})=\frac{1}{2}\log[(2\pi e)^{\dim Z}|\bm{\Sigma}_{M}|]$ [35]. Substituting into Eqn 20, we have with probability at least $1-\delta$,

### v3

@@ 3.2 Few-Shot Generalization to In-Distribution Data
For each environment, the meta-policy is trained on 20 random tasks. To test IID generalization, we sample one trajectory for each testing task as the context, which enables few-shot adaptation of the agent via an updated task representation $\bm{z}$. Fig. 4 illustrates the learning curves of UNICORN vs. all baselines in terms of testing returns, which correspond to the IID entries of Table 2. The results demonstrate that UNICORN variants, especially UNICORN-SS, consistently achieve SOTA performance across all benchmarks. The observation that the asymptotic performance of CSRO is comparable to UNICORN is expected since it also optimizes a linear combination of the lower and upper bound of $I(\bm{Z};\bm{M})$ by Theorem 2.3, with a different implementation choice. Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process. Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired. Moreover, the average training time for MACAW is about three times longer than the context-based methods.
@@ 3.3 Few-Shot Generalization to Out-of-Distribution Behavior Policies
Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:n_{M},t=0:n}$ the checkpoints of behavior policies for all tasks logged during data collection, where $i$ refers to the task index and $t$ represents training iteration. To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:n_{M},t=0:n}}$ to collect the OOD contexts and the conditioned rollouts. The OOD testing performance is measured by averaging returns across all testing tasks and behavior policies, as shown in the OOD entries of Table 2. While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a significant margin. Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT. This also justifies the storyline of COMRL development from 2021 to date
@@ 3.4 Influence of Data Quality
Table 3: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged across 6 random seeds. The best is bolded and the second best is underlined.
@@
We observe that CORRO fails on all three datasets, which is likely attributable to the reliance of CORRO on the negative sample generator. On one hand, under narrow data distribution, there might be little overlap among high-density state-action supports of different tasks. On the other hand, samples collected by random policies are often indistinguishable across tasks [22]. Both factors make it more challenging to train a robust generator for producing high quality negative samples conditioning on specific state-action pairs. As for Prompt-DT, its performance improves significantly with increased data quality, which is expected since the decision transformer adopts a behavior-cloning-like supervised learning style.
@@ 4.1 Is UNICORN Model-Agnostic?
As UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic. Hence a natural idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
We employ a straightforward implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT, akin to in-context learning for large language models [45] and in-context RL [46, 47, 48, 49]. Instead of the non-trainable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$. We name these variants UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
As shown in Table 4 and Fig. 7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option. Despite the gap between our implementation of UNICORN-DT and its MLP counterpart UNICORN-SS in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [50] and DT [51]. We leave this verification to future work.
@@ 4.2 Can UNICORN be Exploited for Model-Based Paradigms?
Figure 5: Testing returns for OOD tasks. The learning curves are averaged over 6 random seeds.
With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ as a world model [52], UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks. We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Fig. 1). Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$. The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent. We adopt the conventional ensemble technique in model-based RL  [53, 54] to stabilize the training process.
As shown in Fig. 5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance (indicating positive few-shot transfer) in this extremely challenging scenario where context shift is induced by OOD tasks.
@@ 5 Conclusion & Limitation
In this paper, we unify three signature context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective. Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift. Based on the proposed framework, we instantiate two novel algorithms called UNICORN-SUP and UNICORN-SS which are demonstrated to be remarkably robust and broadly applicable through extensive experiments. We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL. Moreover, we expect our framework to incorporate almost all existing COMRL methods that tackle task representation learning. Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic $\bm{Z}$ of $\bm{X}$ w.r.t. $\bm{M}$, it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound approximation like the ones introduced by  Theorem 2.3, up to some regularizations or constraints. With such mathematical generalization as well as the empirical evidence of UNICORN being model-agnostic (see Section 4.1), we envision it as a nascent offline pre-training paradigm of foundation models for decision making [21].
Limitation  Since our framework assumes a decoupling of task representation learning and offline policy optimization, it does not directly elucidate how high-quality representations are indicative of higher downstream RL performance. Another limitation would be the scale of the experiments. Our offline datasets cover at most 40 tasks and $\sim 450k$ transition tuples, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small to reduce the approximation error in Theorem 2.4). However, we expect our conclusion to extrapolate to larger task sets, dataset sizes and model parameters, for which the DT variants may demonstrate better scaling. Lastly, in principle, the information theoretic formalization in this work should be directly applicable to online RL. However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B for details), which are apparently violated in the online scenario. Extending our framework to online RL is interesting and nontrivial, which we leave for future work.
@@ Appendix B Proofs
Remark: Consider a practical implementation of $p(\bm{z}|M)$ as a multivariate Gaussian $\mathcal{N}(\bm{\mu}_{M},\bm{\Sigma}_{M})$, which gives $H(\bm{Z}|\bm{M})=\frac{1}{2}\log[(2\pi e)^{\dim Z}|\bm{\Sigma}_{M}|]$ [32]. Substituting into Eqn 21, we have with probability at least $1-\delta$,